resistance of a wire is 10.5 W at 21degree celsius and at 147 degree celsius ut is 16.5 W . Find its temperature coefficient of resistance

Grade:12

1 Answers

Eshan
5 years ago
Dear student,

Resistance of wire after increasing the temperature becomes

R=R_0(1+\alpha\Delta T)
\implies 16.5=10.5(1+\alpha(126^{\circ}C))
\implies \alpha=4.5\times 10^{-3}/^{\circ}C